# Preconditioners for the Stochastic Training of Implicit Neural Representations

### v1

@@ 3.3 Why INRs train so well with Adam compared to SGD
Given an objective function $L(\theta)=|f(\theta)|^{2}$ for a non-linear function $f$. The Jacobian $J_{L}$ of $L$ can be computed via the chain rule: $J_{L}=2J_{f}^{T}f$, where $J_{f}$ denotes the Jacobian of $f$. The Hessian of $L$, $H_{L}$, can be computed by one more application of the chain rule
@@
We can then approximate the Hessian $H_{L}$ by the term $2J_{f}^{T}J_{f}$. This leads to
@@
Eqn. 13 yields the basic idea of the Gauss-Newton method. By taking the Gauss-Newton matrix $J_{L}^{T}J_{L}$ as an approximation to the Hessian $H_{L}$, we can perform an optimization update that uses the step: $\Delta_{GN}=-\frac{1}{J^{T}J}\nabla L$. The Gauss-Newton matrix is in general only positive semi-definite. Therefore, when implementing the algorithm one often adds a small damping factor $\lambda$ so the implemented step is $\Delta_{GN}=-\frac{1}{J^{T}J+\lambda}\nabla L$. A similar derivation can be made for more general loss functions used in non-linear least squares type regression problems [37]. It follows that the Gauss-Newton method can be seen as a preconditioned SGD that uses a Hessian approximate preconditioner given by the Gauss-Newton matrix $J^{T}J$.
The Adam optimizer involves two key steps. The first is to compute a first moment estimate of the form $m_{t}=\beta_{1}m_{t-1}+(1-\beta_{1})g_{t}$, where $g_{t}$ is the gradient of a given objective function $f$ at iteration $t$, and $m_{0}=0$. The second is to compute a second moment estimate $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})g_{t}^{2}$, where $g_{t}^{2}=g_{t}\odot g_{t}$, $\odot$ denotes the Hadamard product, and $v_{0}=0$. The numbers $10<\beta_{1}<\beta_{2}<$ 1 are hyperparameters. The algorithm then applies a bias correction to the above first and second moment estimates, and then takes a moving average to perform an update step based on $-\frac{m_{t}}{\sqrt{v_{t}}}$, see [26] for details.
@@
where $J_{t}$ denotes the Jacobian at iteration $t$. Thus we see that Adam applies a preconditioner given by the diagonal of the Gauss-Newton matrix. Since the Gauss-Newton matrix can be seen as a first order approximation to the Hessian, Adam can be seen as using curvature information to adapt the learning rate for faster optimization. This suggests that Adam can outperform SGD which employs no preconditioner.
We empirically verified this insight in Fig. 4 on a binary occupancy and NeRF experiment. The details of the architectures and experiments are given in Sec. 4.2 and 4.3. It is clear from this figure that Adam outperformed SGD in both signal modalities. Similar results hold in the case of a ReLU/ReLU-PE INR, see supp. material.
3.4 Activation matters for curvature preconditioning
In Sec. 3.1, we highlighted that many implemented preconditioners in practice do not directly compute a Hessian matrix. Instead, they rely on approximations through Hessian vector products. Consequently, comprehending the Hessian vector product’s structure becomes crucial for INRs.
@@ 3.4 Activation matters for curvature preconditioning
Let $F$ denote a sine- or Gaussian-activated INR. Let $\mathcal{L}(\theta)$ denote the MSE loss associated to $F$ and a training set $(X,Y)$. Let $H$ denote the Hessian of $\mathcal{L}$ at a fixed parameter point $\theta$. Then given a non-zero vector $v$ the Hessian vector product $Hv$ will be a dense vector.
@@
Let $F$ denote a ReLU/ReLU-PE-activated INR. Let $\mathcal{L}(\theta)$ denote the MSE loss associated to $F$ and a training set $(X,Y)$. Let $H$ denote the Hessian of $\mathcal{L}$ at a fixed parameter point $\theta$. Then given a non-zero vector $v$ the Hessian vector product $Hv$ will be a sparse vector.
Theorems 3.1 and 3.2 also hold for the binary cross-entropy loss.
Theorem 3.2 implies that when applying a preconditioner, built from a Hessian vector product, to a ReLU/ReLU-PE INR, such as ESGD [16] or AdaHessian [66], only a few components of the gradient will be scaled during the update step. This means that the preconditioner will not be adding much to the optimization process. On the other hand, theorem 3.1 implies that applying a curvature preconditioner to the gradient of the MSE loss of a sine/Gaussian-activated INR should help optimization as many of the gradient components will get scaled by a quantity depending on the curvature of the loss landscape.
Fig. 3 shows the results of a Hessian vector product of the MSE loss of a ReLU-PE and Gaussian INR with respect to the weights in the penultimate layer, see Sec. 4.1 for the network configuration. The random vector was chosen from a Rademacher distribution. Hessian vector product with respect to weights in the other layers of the neural network exhibited similar properties. As can be seen from the experiment the ReLU-PE network produces a sparse vector, as implied by theorem 3.2, and the Gaussian one produces a dense vector, as implied by theorem 3.1.
Theorem 3.1 suggests that training a ReLU/ReLU-PE INR with a curvature aware preconditioner that uses Hessian vector products would not lead to any significant advantage. In order to empirically verify this, we trained a ReLU-PE INR on a NeRF task, see Sec. 4.3 for details on the configuration, using ESGD and Adam. As shown in Fig. 2, the ReLU-PE INR trained with Adam significantly outperforms the one trained with ESGD. This suggests that for ReLU/ReLU-PE activated INRs, such curvature aware preconditioners offer no benefit.
Figure 4: Comparison of training convergence between Adam and SGD on optimizing a 3D occupancy field (Left) and a NeRF (Right). Binary cross-entropy (BCE) loss measures the accuracy of the shape reconstruction, while PSNR loss assesses the scene fidelity.
Figure 5: Comparison of training convergence in Mean Squared Error (MSE) loss (a) and computational complexity of various preconditioner (b and c). ESGD demonstrates superior performance compared to other curvature-aware preconditioners on an image reconstruction task, striking a balance between accuracy and computational efficiency.
@@ 4 Experiments
We compare different preconditioners in training of INR on various popular benchmarks [25, 55, 54, 31]: 2D image reconstruction, 3D shape reconstruction and novel view synthesis using NeRF. We focus on analyzing Gaussian networks, with variance denoted by $\sigma$, with preconditioners, which are ESGD [16] [^1] https://github.com/crowsonkb/esgd, AdaHessian [66] [^2] https://github.com/amirgholami/adahessian, Shampoo [24] [^3] https://github.com/google-research/google-research/tree/master/scalable_shampoo/pytorch, and Kronecker-based preconditioner [33] [^4] https://github.com/Thrandis/EKFAC-pytorch, see supp. material for analysis for other non-traditional activations and the reproducibility details. Remark: We observed that in the implementation of AdaHessian, an equilbrated preconditioner $D^{E}$ is utilized, even though the paper primarily focuses on using a Jacobi preconditioner [16]. Therefore, we introduced a Jacobi preconditioner $D^{J}$ into AdaHessian, which we refer to as AdaHessian(J). We use the term AdaHessian(E) for the original implementation. Since the local Hessian computed with minibatches is generally noisy [66], we opted to recompute the $D^{E}$ for every $N$ iterations for computational efficiency, we use $N=50$ for all the experiments. Additionally, we also employed a moving average technique [66, 26] to mitigate the impact of noise.
@@ 4.1 Image Reconstruction
Given pixel coordinates $\mathbf{x}\in\mathbb{R}^{2}$, we optimized an INR to regress the associated RGB values $c\in\mathbb{R}^{3}$. We use a 5-layer Gaussian-activated network with 256 neurons and $\sigma=0.05$. The network is trained stochastically using mini-batches of size $512$. As seen in Fig. 5(a), while both AdaHessian(J) and AdaHessian(E) demonstrate faster convergence than ESGD in terms of iterations (Fig. 5b), it is important to note that their training time per iteration is longer than ESGD.
4.2 3D Shape Reconstruction
Given 3D point clouds $\mathbf{x}\in\mathbb{R}^{3}$, we optimize an INR to represent the occupancy field of 3D shapes.  [^5] http://graphics.stanford.edu/data/3Dscanrep/ For training, we sampled 12 million 3D points – one-third of the points were sampled uniformly within the volume, and remaining two-thirds of the points were sampled near the mesh surface and perturbed with random Gaussian noise using sigma of 0.1 and 0.01, respectively [25]. We used a 5-layer Gaussian network with 256 neurons and $\sigma=0.09$. The network was trained stochastically using mini-batches of size $50000$. We optimized a binary cross-entropy loss between the predicted occupancy and the true occupancy. Fig. 6(a) shows that ESGD has superior convergence than other preconditioner-type optimizers, see more results in supp. material. Interestingly, when dealing with higher mode signals, AdaHessian(J) and AdaHessian(E) no longer showed superior convergence, as in Sec. 4.1. We speculated that the local Hessians might become noisy in higher mode signals, thereby impacted the gradient updates. However, this noise did not have a significant impact on ESGD’s convergence since the preconditioner updates were performed every $N$ iterations.
Figure 6: Comparison of training convergence (a) in average of Binary Cross- Entropy (BCE) loss of various preconditioners on optimizing a 3D occupancy field. We show the intermediate results (b) at epoch 500 for Adam [26](Left) and ESGD [16](Right). Compared to INR trained with Adam, ESGD has reconstructed the shapes with significantly improved fidelities (zoomed in for better visibility).
4.3 NeRF
NeRF is a compelling strategy for utilizing an INR to model 3D objects and scenes using multi-view 2D images. This approach shows promise for generating high-fidelity reconstruction in novel view synthesis task [35, 47, 15, 29]. Given 3D points $\mathbf{x}\in\mathbb{R}^{3}$ and viewing direction, NeRF estimates the radiance field of a 3D scene which maps each input 3D coordinate to its corresponding volume density $\sigma\in\mathbb{R}$ and directional emitted color $\mathbf{c}\in\mathbb{R}^{3}$ [35, 29, 15]. We used a 8-layer Gaussian-activated network with 256 neurons and $\sigma=0.1$. Table 1 shows that ESGD achieves on average faster convergence than Adam, see more results in the supp. material.
@@ 4.3 NeRF
Table 1: Quantitative results of NeRF on instances from the LLFF dataset [34]. Note that the report time solely refers to the optimization update step.
@@ 5 Conclusion
We investigated the use of preconditioners in stochastic training of INRs. We found that employing a curvature-aware preconditioner significantly speeds up training for INRs admitting non-traditional activations such as Gaussian functions. Experimental validation across signal modalities, including images, shape reconstruction, and neural radiance fields, using ESGD and AdaHessian algorithms supported our findings. However, for INRs with ReLU/ReLU-PE activations, curvature-aware preconditioners offered no significant advantage over Adam. This was confirmed in a comparison study on a NeRF task, where Adam outperformed ESGD for a ReLU-PE INR.
6 Limitations
Our investigation into condition numbers, as detailed in Sec. 3.2, revealed that an equilibrated preconditioner effectively reduces the condition number of the loss landscape compared to a Jacobi preconditioner. This theoretical insight was established in [16] and further validated through empirical experiments involving autoencoder tasks. Surprisingly, our image regression experiment (see Figure 5) demonstrated that a Jacobi preconditioner yielded faster training results than an equilibrated preconditioner. These findings suggest that when employing a preconditioned gradient descent algorithm for optimizing implicit neural representations, the influence of the condition number on convergence may be more complex than initially anticipated. Further research in this area is warranted to gain a deeper understanding.

### v2

@@ 4.3 Training with Adam
Given an objective function $L(\theta)=|f(\theta)|^{2}$ for a non-linear function $f$. The Jacobian $J_{L}$ of $L$ can be computed via the chain rule: $J_{L}=2J_{f}^{T}f$, where $J_{f}$ denotes the Jacobian of $f$. The Hessian of $L$, $H_{L}$, can be computed by an application of the chain rule
@@
We can then approximate the Hessian $H_{L}$ by the term $2J_{f}^{T}J_{f}$. Leading to the first order approximation
@@
Eq. 9 yields the basic idea of the Gauss-Newton method. By taking the Gauss-Newton matrix $J_{L}^{T}J_{L}$ as an approximation to the Hessian $H_{L}$, we can perform an optimization update that uses the step: $\Delta_{GN}=-\frac{1}{J^{T}J}\nabla L$. The Gauss-Newton matrix is in general only positive semi-definite. Therefore, when implementing the algorithm one often adds a small damping factor $\lambda$ so the implemented step is $\Delta_{GN}=-\frac{1}{J^{T}J+\lambda}\nabla L$. A similar derivation can be made for more general loss functions used in non-linear least squares type regression problems [34]. It follows that the Gauss-Newton method can be seen as a preconditioned SGD that uses a first order (first derivatives) Hessian approximate preconditioner given by the Gauss-Newton matrix $J^{T}J$.
Using the above observations we now show how to view Adam as an optimizer that applies a diagonal preconditioner. The Adam optimizer involves two key steps. The first is to compute a first moment estimate of the form $m_{t}=\beta_{1}m_{t-1}+(1-\beta_{1})g_{t}$, where $g_{t}$ is the gradient of a given objective function $f$ at iteration $t$, and $m_{0}=0$. The second is to compute a second moment estimate $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})g_{t}^{2}$, where $g_{t}^{2}=g_{t}\odot g_{t}$, $\odot$ denotes the Hadamard product, and $v_{0}=0$. The numbers $10<\beta_{1}<\beta_{2}<$ 1 are hyperparameters. The algorithm then applies a bias correction to the above first and second moment estimates, and then takes a moving average to perform an update step based on $-\frac{m_{t}}{\sqrt{v_{t}}}$, see [23] for details.
@@
where $J_{t}$ denotes the Jacobian at iteration $t$. Thus we see that Adam applies a preconditioner given by the diagonal of the Gauss-Newton matrix. Since the Gauss-Newton matrix can be seen as a first order approximation to the Hessian, Adam can be seen as using an approximation to curvature to adapt the learning rate for faster optimization. This suggests that Adam can outperform SGD which employs no preconditioner.
We empirically validated this insight across three distinct neural field tasks: image reconstruction, binary occupancy, and NeRF. For each task, we compared the performance of ReLU(PE) activated networks trained with both SGD and Adam against networks using wavelet, sine, and Gaussian activations. As shown in Fig. 3, Adam consistently outperforms SGD in all cases. This suggests that Adam’s use of a preconditioner significantly enhances its performance for neural field applications, see Sec. 5 for the details of the architectures and experimental setup.
4.4 Main theorems
Figure 4: Percentage sparsity of Hessian-vector product matrices across all layers throughout the training process of a 2D image reconstruction task, using ESGD to train networks with different activations. The x-axis ranges from 00 to $1$, with values closer to 1 indicating higher sparsity.
In Sec. 4.3, we discussed how Adam can be viewed as applying a diagonal preconditioner based on the diagonal of the Gauss-Newton matrix, which uses first derivatives to approximate the Hessian matrix. Here, we aim to identify the architectural properties that enable effective training of neural fields with diagonal preconditioners that give a better approximation to the Hessian. Our primary theorems demonstrate that the choice of activation function is crucial for enabling efficient training with such preconditioners, highlighting its central role in achieving effective curvature-aware optimization for neural fields.
Recall Sec. 4.1 that diagonal preconditoners such as the equilibrated precondition $D^{E}$ or the Jacobi preconditioner $D^{J}$ all involve the Hessian matrix. In the context of neural networks the Hessian of the loss landscape is often an extremely large matrix. Thus in practical settings one often replaces the Hessian with a Hessian vector product which is much cheaper to store [34].
Our first theorem shows that the Hessian vector product of the loss landscape associated to a sine, Gaussian, wavelet or sinc activated neural field produces a dense vector.
@@ 4.4 Main theorems
Let $F$ denote a sine, Gaussian, wavelet or sinc activated neural field. Let $\mathcal{L}(\theta)$ denote the MSE loss associated to $F$ and a training set $(X,Y)$. Let $H$ denote the Hessian of $\mathcal{L}$ at a fixed parameter point $\theta$. Then given a non-zero vector $v$ the Hessian vector product $Hv$ will be a dense vector.
@@
Let $F$ denote a ReLU/ReLU(PE)-activated neural field. Let $\mathcal{L}(\theta)$ denote the MSE loss associated to $F$ and a training set $(X,Y)$. Let $H$ denote the Hessian of $\mathcal{L}$ at a fixed parameter point $\theta$. Then given a non-zero vector $v$ the Hessian vector product $Hv$ will be a sparse vector.
Theorem 4.2 and Theorem 4.3 also hold for the binary cross-entropy loss. The proofs of Theorem 4.2 and Theorem 4.3 are given in the supp.
Theorem 4.3 implies that when applying a preconditioner, built from a Hessian vector product, to a ReLU/ReLU(PE) INR, such as ESGD [14] or AdaHessian [62], only a few components of the gradient will be scaled during the update step. This means that the preconditioner will not be adding much to the optimization process. On the other hand, Theorem 4.2 implies that applying a preconditioner built from the Hessian to the gradient of the MSE loss of a sine, Gaussian and wavelet-activated neural field should help optimization as many of the gradient components will get scaled by a quantity depending on the curvature of the loss landscape. This yields a clear difference between ReLU/ReLU(PE) activated neural fields and those activated with a sine, Gaussian, wavelet activation. In particular, Theorem 4.2 implies that neural fields admitting a sine, Gaussian, wavelet activation have the ability to be trained efficiently with a diagonal curvature aware preconditioner such as the equilibrated preconditioner $D^{E}$ defined in Sec. 4.1.
Fig. 4 shows the results of a Hessian vector product of the MSE loss of a ReLU(PE) and Gaussian neural field for an image reconstruction task for the weights in all layers of the network layer throughout training. The random vector was chosen from a Rademacher distribution and the reported figures are given as the mean of the Hessian of all layers averaged over all iterations of training. As demonstrated by the experiment, the ReLU(PE) network generates a sparse Hessian-vector product, as predicted by Theorem 4.3, while the Gaussian network produces a dense vector, as indicated by Theorem 4.2. Furthermore, as we will observe in Sec. 5, this sparsity helps explain why Adam performs well for neural fields with ReLU (ReLU(PE)) activations, whereas an algorithm like ESGD [14] does not. This is because Adam does not rely on Hessian-vector product.
@@ 5 Experiments
Figure 5: Comparison of training convergence and computational complexity for various preconditioners. We evaluated a Gaussian-activated neural field on the lion instance from the DIV2K dataset. ESGD demonstrates superior convergence compared to other preconditioners, striking a balance between accuracy and computational efficiency. Note: We provide the time-based comparisons in the supp. (Sec 3.1). Similar analysis for other activations and additional instances from the DIV2K dataset are available in the supp. (Sec. 3.2).
We compared various preconditioners over popular benchmarks [22, 54, 50, 28, 52]: 2D image reconstruction, 3D shape reconstruction and novel view synthesis with NeRF. Our primary focus was on curvature-aware preconditoners that are well-suited for the stochastic training of neural network. Specifically, we compared the following methods: ESGD [14] [^2] https://github.com/crowsonkb/esgd, AdaHessian [62] [^3] https://github.com/amirgholami/adahessian, Shampoo [21] [^4] https://github.com/google-research/google-research/tree/master/scalable_shampoo/pytorch, and Kronecker-based preconditioner [30] [^5] https://github.com/Thrandis/EKFAC-pytorch. We provide the details of each algorithm including the reproducibility details of each experiment and machines we used to run each experiment in the supp. In the main paper, we focused on analyzing Gaussian networks, see supp. for additional analysis for other non-traditional activations and other neural field tasks.
Remark for AdaHessian
We observed that in the implementation of AdaHessian, an equilbrated preconditioner $D^{E}$ is utilized, even though the paper primarily focuses on using a Jacobi preconditioner [14]. Therefore, we introduced a Jacobi preconditioner $D^{J}$ into AdaHessian, which we refer to as AdaHessian(J). We use the term AdaHessian(E) for the original implementation.
Remark for ESGD
We further modified the ESGD by applying the squared Hessian diagonal preconditioner in two settings: using an exponential moving average (ESGD) and using an inifinity norm (ESGD-max) [23]. We also opted to recompute the $D^{E}$ for every $N$ iterations for computational efficiency, we use $N=100$ for all the experiments.
@@ 5.1 Image Reconstruction
We evaluated on a 2D image reconstruction task, where given pixel coordinates $\mathbf{x}\in\mathbb{R}^{2}$, we optimized a neural network to regress the corresponding RGB values $c\in\mathbb{R}^{3}$. We used a 5-layer Gaussian-activated network with 256 neurons. We used $\sigma=0.05$ for the Gaussian activation. We trained the network using mini-batches of size $512$. We compared the training convergence of different preconditioners and the Adam optimizer on the lion instance from the DIV2K dataset [1] in Fig. 5(a). Additional results on the rest of the DIV2K dataset are provided in the supp. As predicted by Theorem 4.2, Gaussian network equipped with the dense Hessian-vector product exhibits faster convergence when optimized with curvature-aware preconditioned methods compared to Adam. Notably, ESGD demonstrates faster convergence and has comparable training time and memory usage per iteration compared to Adam.
ESGD and AdaHessian obtain Hessian information by computing Hessian-vector products, which makes them very computational efficient. Each Hessian-vector product is generally as fast as a single gradient computation, as demonstrated in Fig. 5(b). In contrast, other preconditioners, particularly Shampoo are more computationally expensive because they involve Kronecker product computations for preconditioner matrices, see supp. for details.
@@
trex 25.50 80K 164.96 22.04 0.71 0.31 120K 209.25 22.21 0.75 0.23
Table 1: Quantitative results of NeRF on all instances from the LLFF dataset [31]. The reported time solely refers to the optimization update step. Note: Results for the Blender synthetic [32] and BLEFF [60] datasets are available in supp.
5.2 3D Binary Occupancy Reconstruction
We evaluated on learning a binary occupancy field, where given 3D point coordinates $\mathbf{x}\in\mathbb{R}^{3}$, we optimized a neural network to represent the occupancy field of 3D shapes. Following existing works [22], we sampled one million 3D points for training – one-third of the points were sampled uniformly within the volume, and the remaining two-thirds of the points were sampled near the mesh surface and perturbed with random Gaussian noise using sigma of 0.1 and 0.01, respectively [22]. We trained the network using binary cross-entropy (BCE) loss, which compares the predicted occupancy and the true occupancy. We used a 5-layer Gaussian network with 256 neurons and $\sigma=0.09$. We trained the network using mini-batches of size $50000$.
We compared the training convergence of different preconditioners and the Adam optimizer on the armadillo instance from the Stanford 3D Scanning dataset.  [^6] http://graphics.stanford.edu/data/3Dscanrep/ Fig. 6 shows that ESGD exhibits superior convergence than other preconditioner-type optimizers. We showcase a qualitative result comparing the mesh reconstruction of ESGD and Adam given the same number of iteration in Fig. 7. Interestingly, when dealing with higher mode signals, both AdaHessian and Shampoo no longer demonstrated superior convergence, unlike in the 2D image reconstruction task in Fig. 5(a). We speculate that the local Hessians may become noisy with higher mode signals, thereby impacting the gradient updates. In contrast, this noise did not significantly affect ESGD’s convergence because its preconditioner updates were performed every $N$ iterations. Hence, we will focus on comparing Adam and ESGD for the experiments in the following section.
Figure 6: Comparison of training convergence for various preconditioners. We evaluated a Gaussian-activated neural field on a 3D binary occupancy reconstruction task on the armadillo instance. ESGD demonstrates superior convergence compared to other preconditioners, striking a balance between accuracy and computational efficiency. Note: Similar analysis for other activations are available in the supp.
Figure 7: Qualitative results for the dragon instance from Stanford 3D Scanning dataset. Compared to the Gaussian-based neural field trained with Adam (Fig. 7(a)), ESGD (Fig. 7(b)) has reconstructed the shapes with significantly improved fidelity at epoch 500 (zoom in 4×4$\times$4 × for better visibility). Note: Qualitative results for the rest of the instances from Stanford 3D Scanning Dataset are available in supp.
5.3 Neural Radiance Fields (NeRF)
Neural Radiance Fields (NeRF) have recently emerged as a pioneering approach in novel view synthesis, utilizing neural networks to model volumetric fields of 3D scenes using multi-view 2D images [32, 44, 12, 26, 13]. Specifically, given 3D points $\mathbf{x}\in\mathbb{R}^{3}$ and viewing direction, NeRF estimates the 5D radiance field of a 3D scene which maps each input 3D coordinate to its corresponding volume density $\sigma\in\mathbb{R}$ and directional emitted color $\mathbf{c}\in\mathbb{R}^{3}$ [32, 26, 12]. We used an 8-layer Gaussian-activated network with 256 neurons and $\sigma=0.1$. Recall that in the previous section we demonstrated that ESGD exhibited superior performance in terms of convergence and computational complexity compared to other preconditioners. Therefore, we focused on comparing Adam and ESGD here. We determined the same training PSNR at which both model converged and reported the corresponding optimization metrics, including the number of iterations and the total time taken for the update computation. We evaluated the PSNR on a hold-out test scene. Table 1. shows that ESGD achieves on average faster convergence than Adam, see more results in the supp.
6 Limitations
While our work theoretically established the suitability of curvature-aware preconditioners for training neural fields with non-traditional activations such as sine, Gaussian, and wavelets, it also highlighted their ineffectiveness for neural fields employing ReLU-PE activations. Specifically, we showed that Adam performs well for ReLU-PE due to its use of a diagonal preconditioner derived from the Gauss-Newton matrix, rather than relying on Hessian vector products. However, this raises an important limitation of our work: we did not explore whether alternative preconditioners, potentially more effective than the Gauss-Newton matrix, could be developed to improve the training of ReLU-PE activated neural fields beyond what Adam achieves. This is particularly relevant since ReLU-PE remains widely used in practical applications. This question represents an avenue for future research, which we leave it as future work.
@@ 7 Conclusion
We explored the impact of preconditioners on the stochastic training of neural fields, developing a theoretical framework based on Hessian-vector products. Our analysis demonstrates that curvature-aware preconditioners can significantly accelerate training for neural fields with non-traditional activations, such as sine, Gaussian, and wavelet. Experimental validation across various tasks—including image reconstruction, shape modeling, and neural radiance fields—using ESGD and AdaHessian confirmed these benefits. In contrast, for neural fields with ReLU or ReLU-PE activations, our theory predicted no significant advantage from curvature-aware preconditioning, highlighting the activation-dependent nature of these methods. For practitioners employing activations like sine, Gaussian, or wavelet, robust curvature-aware algorithms provide an efficient approach for stochastic training across diverse applications.